Consider an economy with two goods produces, computers and bicycles. The following Table summarizes the production and price information of these two goods over two years: Price Quantity

Year 1

computers: price:500 quantity:10

bicycles: price:50 quantity:100

Year 2

computers: price:600 quantity:20

bicycles: price:80 quantity:120

2.1 Compute the Nominal GDP in both years. What is the rate of growth of nominal GDP?

2.2 Compute the Real GDP in both years using Year 1 as the base year. What is the rate of growth for real GDP?

2.3 Compute the implicit GDP price deflator, and the inflation from year 1 to year 2.

2.4 Suppose that computers in year 2 are of higher quality compared to computers in year 1. How would such information bias your results? Explain.
Answer #1

2.1 Nominal GDP= Price*Quantity

YEAR 1

GDP from computer production= 500*10=$5000

GDP from bicycle production= 50*100= $5000

Total nominal GDP in year 1= $5000+$5000=$10,000

YEAR 2

GDP from computer production= 600*20=$12000

GDP from bicycle production= 80*120= $9600

Total nominal GDP in year 2= $12000+$9600=$21600

2.2 Real GDP= Quantity of given year*Prices of base year

YEAR 1

GDP from computer production= 500*10=$5000

GDP from bicycle production= 50*100= $5000

Total GDP in year 1= $5000+$5000=$10,000

YEAR 2

GDP from computer production= 500*20=$10000

GDP from bicycle production= 50*120= $6000

Total GDP in year 2= $10000+$6000=$16000

% Rate of growth of real GDP= (Real Gdp in year 2- Real Gdp in year1/Real Gdp in year 1)*100

=(16000-10000/10000)*100

=60%

2.3 % GDP Deflator= (Nominal Gdp/ Real Gdp)*100

Year 1

= (10000/10000)*100

=100

Year 2

=(21600/16000)*100

=135%

Inflation rate= (GDP deflator in year 2-GDP deflator in year 1/GDP deflator in year 1)*100

=(135-100/100)*100

=35%

Hence, inflation rate is 35%

2.4 One problem with system of national accounts is that quality is not taken into consideration. A firm may charge higher price due to better quality product it is offering and prices may rise to reflect the same. A price rise or inflation of this kind is not always a problem, rather technological progress is vital for growth of the economy and price rise due to better quality products is not a problem always.

On knowing this information one might wish to understate the price in year two since a portion of the price in year two is reflecting quality progress and is not giving a uniform comparison between the two products in different years.